A large container which has 258 g of water initially at 23 °C is placed in a microwave oven. The magnetron in the microwave oven can produce 950 W of EM radiation. The microwave is turned on for 2.5 minutes.

a) Assuming that all the microwave energy is absorbed by the water, how much water is left in the container?

b) The OXO 8716900 electric kettle (1500 W) manufacturer’s claim that it can boil up to 1.75 L of water faster than the microwave, safer than the stovetop, and runs off a mains voltage of 120 V. Compare the efficiency of the two approaches.

c) Name the different ways of heat loss in these systems (Microwave & Kettle) during the heating process.
1
Expert's answer
2020-01-31T10:43:06-0500

a) According to law of conservation of energy, we can write

"P\u00d7t=Mc(t_2-t_1)+mL (1)"

Using (1) we get

950×2.5×60=0.258×4200×(100-23)+2300000×m

m=25.7 g

Answer

25.7 g


b)

for microwave:

"\frac{950}{1000} =0.95"


For kettle:

"\frac{1500×0.375}{1000} =0.56"


c)

For kettle:

a scum is generated on the heating element, metal walls of the case

all this increases electricity losses

For microwave:

some of the energy goes to heat the air inside